### v1

@@ Abstract
Simulation frameworks play a pivotal role in the safe development of robotic applications. However, often different components of an envisioned robotic system are best simulated in different environments/simulators. This poses a significant challenge in simulating the entire project into a single integrated robotic framework. Specifically, for partially-open or closed-source simulators, often two core limitations arise. i) Actors in the scene other than the designated robots cannot be controlled during runtime via interfaces such as ROS and ii) retrieving real-time state information (such as pose, velocity etc.) of objects in the scene is prevented. In this work, we address these limitations and describe our solution for the use case of integrating aerial drones simulated by the powerful simulator Sphinx (provided by Parrot Drone) into the Gazebo simulator. We achieve this by means of a mirrored instance of a drone that is included into existing Gazebo-based environments. A promising application of our integrated simulation environment is the task of target tracking that is common in aerial multi-robot scenarios. Therefore, to demonstrate the effectiveness our our integrated simulation, we also implement a model predictive controller (MPC) that outperforms the default PID-based controller framework provided with the Parrot’s popular Anafi drone in various dynamic tracking scenarios thus enhancing the utility of the overall system. We test our solution by including the Anafi drone in an existing Gazebo-based simulation and evaluate the performance of the MPC through rigorous testing in simulated and real-world tracking experiments against a customized PID controller baseline. Source code is published on https://github.com/robot-perception-group/anafi_sim.
Keywords:
Simulation Aerial robotics Real world experiments.
1 Introduction
@@
It is common in research to design and build customized robots for specific purposes and experiments, which is usually a time-consuming process until all hardware-related issues have been solved. In this regard, commercial robotic platforms often provide a shortcut to quick deployment of a robotic system since these issues have been ruled-out by the manufacturer during the development phase of the product. However, for scientific robotic applications this advantage is often diminished when no or only a simulation with limited interfaces or customization options is provided along with the robot.
@@
In this work we solve the aforementioned accessibility problems while focusing on a powerful consumer drone called Anafi that is manufactured by Parrot Drone [11]. For development, testing and mission planning a comprehensive simulation environment named Sphinx [12] is provided. Even-though scenes in which the drone operates can be customized using static and dynamic actors, being only partially open-source Sphinx suffers from the two aforementioned accessibility problems. This is a major drawback since the information about objects in the scene other than the Anafi drone is frequently required, e.g. during the design of a tracking controller and therefore the first limitation in using Sphinx.
The second limitation in using Sphinx is the PID-based controller framework provided with the firmware of the Anafi drone. A common application in aerial robotics is target tracking, e.g. [26, 25]. However, depending on the manoeuvrability of the target, agile tracking controllers are required. Feedback controllers such as the provided PID-based control framework are suitable only for slowly moving targets as otherwise large tracking offsets can occur. Furthermore, the PID control stack integrated in the firmware of the Anafi drone shows a peculiar behaviour considerably limiting its applicability to target tracking tasks. Frequent updates to the target waypoint, a common practice in tracking scenarios, do not yield the anticipated smooth transition between successive waypoints of the drone. Instead, an oscillating motion characteristic can be observed, leading to slow response behaviour which is illustrated by Figure 2.
@@
We demonstrate the efficacy of our two solution approaches as follows. First, we include the Anafi drone in the airship simulation [22]. Second, we evaluate the performance of the MPC in simulated and real-world experiments by comparing it to our customized PID controller baseline. The remainder of this paper is structured as follows. Related work is summarized in Section 2, Section 3 contains information about the connection between Sphinx and Gazebo as well as the controller designs, followed by Section 4 explaining details of the implementation. The results are described in Section 5 before we conclude this work in Section 6.
2 Related Work
@@
Sphinx [12] is a simulator specifically designed for the simulation of the Anafi drone series distributed by Parrot Drone. Whereas both PX4-SITL and RotorS are fully open source simulators, the Anafi drone firmware and Sphinx simulator are proprietary, although customizable to a certain degree. The big advantage in using Sphinx is that it provides a testbed for scenarios involving the high-end drones of the Anafi series. Interfacing the drone firmware and thus retrieving the state of the drone is possible using the Olympe API [6] for which anafi_ros [24] is a ROS-wrapper. This enables easy integration into ROS frameworks, however, integrating the Anafi drones into arbitrary Gazebo-based environments is not possible. anafi_autonomy [1] is build on top of anafi_ros and implements among mission planning features also a velocity controller. In order to also provide a position and thus a tracking controller, we not only implement our own PID-based tracking framework but also an MPC which builds on top of [26]. In this work, an efficient formulation of a MPC for tracking tasks featuring obstacle avoidance with high execution rate has been introduced. We augment the dynamic model of the MPC with additional cummulative error states to minimize the constant tracking error while leaving out unnecessary components such as obstacle avoidance. For this task, we rely higher-level trajectory planning methods that send waypoints to the MPC.
3 Methodology
3.1 Sphinx Simulator
The drones provided by Parrot Drone are known for having long flight times and flight range and are equipped with powerful hardware and sensors for different applications as illustrated in [24], thus making them attractive also for scientific purposes. The publicly available simulator Sphinx [12] allows the simulation of the drones Anafi [8] and Anafi Ai [9] in customizable environments. Sphinx’ physics back-end is based on a customized version of Gazebo 11 and the visualization front-end on a customized version of the graphic engine Unreal Engine 4.26 [10]. However, one major drawback of Sphinx is that it cannot simulate more than one drone instance at once [17].
@@ 3.1 Sphinx Simulator
Static actors, i.e. static meshes, can be placed in the Sphinx environment in customizable poses. Furthermore, Parrot provides a set of ready-to-use dynamic actors such as cars and pedestrians for which trajectories can be defined before runtime [18] using configuration files. User generated actors with customized appearance and motion characteristics can be realized in the modified UE4 editor provided by Parrot Drone. For this purpose Blueprint functions can be used, a graphical programming approach resembling in its workflow MATLAB / Simulink. However, as outlined in Section 1, state information of actors in the scene can neither be updated nor retrieved from outside Sphinx. Technically, numerous plugins exist for updating actor states from outside UE4-based applications, e.g. also via ROS. Retrieving state information from inside the application could also be realized using C++ code which generally can be included in the UE4 editor. However, Parrot Drone has blocked the usage of plugins [14] and custom C++ code [3], necessitating other approaches for multi-robot scenarios.
3.2 Connection between Sphinx and Gazebo
@@
Figure 3: Overview of the dataflow and connection between the components enabling the interface between Sphinx and Gazebo. Simulators are colored orange, components relying on ROS are colored blue. The simulated firmware is a stand alone module provided by Parrot Drone and colored green. The boxes next to the arrows indicate the tool managing the communication with the respective components.
The Sphinx component is an instance of the Sphinx simulator running an empty world in which the Anafi drone is spawned. The drone parameters can be accessed via the simulated firmware implemented as a daemon called Firmwared, as is indicated by the Simulated Firmware block. Parrot Drone provides a Python-based API called Olympe to communicate with the firmware. The component anafi_ros [24] constitutes a ROS wrapper for Olympe which allows sending and retrieving drone information via ROS topics, a necessity to realize other functionalities that are implemented elsewhere in the ROS system, e.g. controllers. The Interface component accesses the Sphinx simulator via a CLI tool provided by Parrot Drone for the purpose of retrieving ground truth data about the Anafi drone. This ground truth data comprises the following. The pose and the linear and angular velocity vectors in both, world frame (an East-North-Up coordinate system) and the body center reference frame. Furthermore, the linear and angular acceleration is available in the body center reference frame only. The gimbal position and orientation and thus the pose of the camera are provided in world frame.
This information is used to set up two objects in Gazebo. The first is a robot model consisting only of a mesh of the Anafi drone for visualization of its pose as retrieved from Sphinx. The second is a camera plugin whose position and orientation is governed by the gimbal of the Anafi drone in Sphinx. Common to both objects is that the influence of gravity is deactivated. This avoids a jittering behaviour in the model’s position which is induced by the influence of gravity in between updates of the pose. Furthermore, collision capabilities are disabled. This is due to the missing possibility to feed collision data back into Sphinx to then have an effect on the flight behaviour of the simulated drone.
A critical issue when running two interfaced simulations in parallel is timing. Both simulations need to run with the same speed as to prevent data loss during exchanging data. Since it is not possible to update Sphinx and Gazebo on a step-wise basis, we address this issue by reducing the real time factor (RTF) to a value which can be reached by both simulations simultaneously. On our setup (see Section 4) this is the case for $RTF=0.6$.
3.3 Controller Design
@@ 3.4 PID Controller Framework
3.5 Quadratic Model Predictive Controller
In order to reduce the tacking error present for PID based tracking controllers, we focus on the reduction of the offset between the moving target waypoint and the drone during the design of the MPC by introducing cumulative error states. In this regard, using a quadratic formulation of the MPC is beneficial as convex quadratic optimization problems are fast to solve and computationally cheap [20].
@@ 3.5 Quadratic Model Predictive Controller
For the required vertical velocity to set the altitude of the drone the value $v_{z}$ taken from $\mathbf{x}_{5}^{*}$ (see (2)) has proven most successful. The yaw angle is not controlled by the MPC but by the same single loop PID controller used in the framework described in Section 3.4.
4 Implementation
4.1 General
@@
  • that our MPC tracking controller (see Section 3.5) outperforms in simulation and real world experiments our customized PID controller framework (see Section 3.4). We introduced the latter to serve as a baseline for the performance evaluation of the MPC while addressing the shortcomings of the PID control stack which is implemented in the Anafi firmware and provided by Parrot Drone.
4.2 Simulation Environment
@@
4.3 Integration of the Anafi Drone into a Multi-Robot Simulation
@@
4.4 Verification of the Custom PID-based Controller Framework and of the MPC
@@
5 Results
5.1 Integration of the Anafi Drone into a Gazebo Simulation
@@
Since Gazebo can have difficulties loading and rendering multiple complex meshes simultaneously, we neglect the display of rotating propellers to keep the model of the Anafi drone as simple as possible.
5.2 Evaluation of Tracking Controllers
General
@@ 5.2.1 Static waypoints
Figure 7 illustrates the same scenario but in the real world experiments. It can be seen that this time the MPC is not beneficial in comparison to the PID controller framework. However, this can be attributed to the windy conditions ($v_{wind}=4$\mathrm{m}\mathrm{/}\mathrm{s}$$ with considerably higher wind gusts possibly present when testing the MPC). Both controllers show slight oscillations when maintaining hover flight indicating the presence of wind gusts.
5.2.2 Straight Line Trajectory
@@ 5.2.3 Circular Trajectory
The MPC shows better performance than the PID across all scenarios, although at higher velocities the advantage of the MPC diminishes. The reason is that in these scenarios the control commands produced by the MPC are bound by the maximum attitude angles of the drone thus limiting its manueverability. The real world experiments confirm the simulation results, although they were conducted under windy conditions subjecting the drone to heavy disturbances. However, the standard deviation of the tracking error $e_{\sigma}$ of the MPC is considerably lower than than the one of the PID controller framework thus indicating that the tracking is managed in an evenly way.
5.2.4 Vertical Rectilinear Periodic Trajectory
@@
6 Conclusion
In this work an approach for using the Sphinx simulator by Parrot Drone to accurately simulate drones within more general Gazebo-based multi-robot simulation frameworks is described. Furthermore, against the background of common tracking tasks, two controller frameworks are introduced to solve limitations present in the existing PID-based controller framework provided by Parrot Drone in the Anafi firmware. Our first controller framework is also PID-based, however, suffers from limited tracking accuracy. For this purpose, our second control framework is an MPC which specifically aims to reduce the tracking error using a modified dynamics model with integral error states. The integration of the Anafi drone is tested using our airship simulation [22] and the controllers are rigorously evaluated in simulation and real world experiments featuring different scenarios.
{credits}
6.0.1 Acknowledgements
The authors thank Curtis Walch and Christian Gall for their support with the flight experiments. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Pascal Goldschmid.
[…]
Both authors’ primary affiliation is the Institute of Flight Mechanics and Controls, University of Stuttgart. Pascal Goldschmid is employed as a research associate and Aamir Ahmad as a tenure-track professor. Aamir Ahmad is also affiliated with the Perceiving Systems department of the Max Planck Institute for Intelligent Systems.

### v2

@@ Abstract
Simulation frameworks are essential for the safe development of robotic applications. However, different components of a robotic system are often best simulated in different environments, making full integration challenging. This is particularly true for partially-open or closed-source simulators, which commonly suffer from two limitations: (i) lack of runtime control over scene actors via interfaces like ROS, and (ii) restricted access to real-time state data (e.g., pose, velocity) of scene objects. In the first part of this work, we address these issues by integrating aerial drones simulated in Parrot’s Sphinx environment (used for Anafi drones) into the Gazebo simulator. Our approach uses a mirrored drone instance embedded within Gazebo environments to bridge the two simulators. One key application is aerial target tracking, a common task in multi-robot systems. However, Parrot’s default PID-based controller lacks the agility needed for tracking fast-moving targets. To overcome this, in the second part of this work we develop a model predictive controller (MPC) that leverages cumulative error states to improve tracking accuracy. Our MPC significantly outperforms the built-in PID controller in dynamic scenarios, increasing the effectiveness of the overall system. We validate our integrated framework by incorporating the Anafi drone into an existing Gazebo-based airship simulation and rigorously test the MPC against a custom PID baseline in both simulated and real-world experiments.
Source code is available at: https://github.com/robot-perception-group/anafi_sim
I Introduction
@@
It is common in research to design and build customized robots for specific purposes and experiments, which is usually a time-consuming process until all hardware-related issues have been solved [3]. In this regard, commercial robotic platforms often provide a shortcut to quick deployment of a robotic system since these issues have been ruled-out by the manufacturer during the development phase of the product. However, for scientific robotic applications this advantage is often diminished when no or only a simulation with limited interfaces or customization options is provided along with the robot.
@@
In aerial robotics, the drones provided by Parrot Drone are known for having long flight times and flight range and are equipped with powerful hardware and sensors for different applications as illustrated in [2], thus making them attractive also for scientific purposes. The publicly available simulator Sphinx [4] allows the simulation of the drones Anafi [5] and Anafi Ai [6] in customizable environments. It is also used by Parrot’s engineers for development and can therefore be considered as a high-fidelity simulation environment for the Anafi drone series. However, even-though scenes in which the drone operates can be customized using static and dynamic actors, being only partially open-source Sphinx suffers from the two aforementioned accessibility problems. This is a major drawback since the information about objects in the scene other than the Anafi drone is frequently required for scientific purposes, e.g. during the design of a tracking controller or in multi-robot simulations. This therefore constitutes the first limitation in using Sphinx we address in this work.
The second limitation of Sphinx we solve in this work is the PID-based controller framework provided with the firmware of the Anafi drone. A common application in aerial robotics is target tracking, e.g. [7, 8]. However, depending on the manoeuvrability of the target, agile tracking controllers are required. Feedback controllers such as the provided PID-based control framework are suitable only for slowly moving targets as otherwise large tracking offsets can occur. Furthermore, the PID control stack integrated in the firmware of the Anafi drone shows a peculiar behaviour considerably limiting its applicability to target tracking tasks. Frequent updates to the target waypoint, a common practice in tracking scenarios, do not yield the anticipated smooth transition between successive waypoints of the drone. Instead, an oscillating motion characteristic can be observed, leading to slow response behaviour which is illustrated by Figure 2.
@@
We demonstrate the efficacy of our two solution approaches as follows. First, we include the Anafi drone in the airship simulation [10]. Second, we evaluate the performance of the MPC in simulated and real-world experiments by comparing it to our customized PID controller baseline. The remainder of this paper is structured as follows. Related work is summarized in Section II, Section III contains information about the connection between Sphinx and Gazebo as well as the controller designs, followed by Section IV explaining details of the implementation. The results are described in Section V before we conclude this work in Section VI.
II Related Work
@@
Sphinx [4] is a simulator specifically designed for the simulation of the Anafi drone series distributed by Parrot Drone. Whereas both PX4-SITL and RotorS are fully open source simulators, the Anafi drone firmware and Sphinx simulator are proprietary, although customizable to a certain degree. The big advantage in using Sphinx is that it provides a testbed for scenarios involving the high-end drones of the Anafi series. Interfacing the drone firmware and thus retrieving the state of the drone is possible using the Olympe API [18] for which anafi_ros [2] is a ROS-wrapper. This enables easy integration into ROS frameworks, however, integrating the Anafi drones into arbitrary Gazebo-based environments is not possible. anafi_autonomy [19] is build on top of anafi_ros and implements among mission planning features also a velocity controller. In order to also provide a position and thus a tracking controller, we not only implement our own PID-based tracking framework but also an MPC which builds on top of [7]. In this work, an efficient formulation of a MPC for tracking tasks featuring obstacle avoidance with high execution rate has been introduced. We augment the dynamic model of the MPC with additional cummulative error states to minimize the constant tracking error while leaving out unnecessary components such as obstacle avoidance. For this task, we rely on higher-level trajectory planning methods that send waypoints to the MPC. Error states in model predictive control are a concept used, for example, in legged robots [20], where the tracking error is defined on a Lie matrix group and linearised in the Lie algebra. [21] builds on this work and extends it for the control of a quadcopter. However, the authors point out shortcomings in the computational speed of their method. The benefit our MPC implementation is that it has even less computational cost than our customized PID controller framework.
III Methodology
III-A Sphinx Simulator
Sphinx’ physics back-end is based on a customized version of Gazebo 11 and the visualization front-end on a customized version of the graphic engine Unreal Engine 4.26 [22]. However, one major drawback of Sphinx is that it cannot simulate more than one drone instance at once [23].
@@ III-A Sphinx Simulator
Static actors, i.e. static meshes, can be placed in the Sphinx environment in customizable poses. Furthermore, Parrot provides a set of ready-to-use dynamic actors such as cars and pedestrians for which trajectories can be defined before runtime [24] using configuration files. User generated actors with customized appearance and motion characteristics can be realized in the modified UE4 editor provided by Parrot Drone. For this purpose Blueprint functions can be used, a graphical programming approach resembling in its workflow MATLAB / Simulink. However, as outlined in Section I, state information of actors in the scene can neither be updated nor retrieved from outside Sphinx. Technically, numerous plugins exist for updating actor states from outside UE4-based applications, e.g. also via ROS. Retrieving state information from inside the application could also be realized using C++ code which generally can be included in the UE4 editor. However, Parrot Drone has blocked the usage of plugins [25] and custom C++ code [26], necessitating other approaches for multi-robot scenarios.
III-B Connection between Sphinx and Gazebo
@@
Figure 3: Overview of the dataflow and connection between the components enabling the interface between Sphinx and Gazebo. Simulators are colored orange, components relying on ROS are colored blue. The simulated firmware is a stand alone module provided by Parrot Drone and colored green. Our customized set of ROS nodes providing the exchange of data between the simulators is colored yellow. The boxes next to the arrows indicate the tool managing the communication with the respective components.
The Sphinx component is an instance of the Sphinx simulator running an empty world in which the Anafi drone is spawned. The drone parameters can be accessed via the simulated firmware implemented as a daemon called Firmwared, as is indicated by the Simulated Firmware block. Parrot Drone provides a Python-based API called Olympe to communicate with the firmware. The component anafi_ros [2] constitutes a ROS wrapper for Olympe which allows sending and retrieving drone information via ROS topics, a necessity to realize other functionalities that are implemented elsewhere in the ROS system, e.g. controllers. For the Interface component we developed a set of ROS nodes accessing the Sphinx simulator via a CLI tool provided by Parrot Drone for the purpose of retrieving ground truth data about the Anafi drone. This ground truth data comprises the following. The pose and the linear and angular velocity vectors in both, world frame (an East-North-Up coordinate system) and the body center reference frame. Furthermore, the linear and angular acceleration is available in the body center reference frame only. The gimbal position and orientation and thus the pose of the camera are provided in world frame.
This information is used by the ROS nodes in the Interface component to set up two objects in Gazebo. The first is a robot model consisting only of a mesh of the Anafi drone for visualization of its pose as retrieved from Sphinx. The second is a camera plugin whose position and orientation is governed by the gimbal of the Anafi drone in Sphinx. Common to both objects is that the influence of gravity is deactivated. This avoids a jittering behaviour in the model’s position which is induced by the influence of gravity in between updates of the pose. Furthermore, collision capabilities are disabled. This is due to the missing possibility to feed collision data back into Sphinx to then have an effect on the flight behaviour of the simulated drone.
A critical issue when running two interfaced simulations in parallel is timing. Both simulations need to run with the same speed as to prevent data loss during exchanging data. Since it is not possible to update Sphinx and Gazebo on a step-wise basis, we address this issue by reducing the real time factor (RTF) to a value which can be reached by both simulations simultaneously.
III-C Controller Design
@@ III-D PID Controller Framework
III-E Quadratic Model Predictive Controller
In order to reduce the tracking error present for PID based tracking controllers, we focus on the reduction of the offset between the moving target waypoint and the drone during the design of the MPC by introducing cumulative error states. In this regard, using a quadratic formulation of the MPC is beneficial as convex quadratic optimization problems are fast to solve and computationally cheap [28].
@@ III-E Quadratic Model Predictive Controller
For the required vertical velocity to set the altitude of the drone the value $v_{z}$ taken from $\mathbf{x}_{5}^{*}$ (see (2)) has proven most successful. The yaw angle is not controlled by the MPC but by the same single loop PID controller used in the framework described in Section III-D.
IV Implementation
IV-A General
@@
  • that our MPC tracking controller (see Section III-E) outperforms in simulation and real world experiments our customized PID controller framework (see Section III-D). We introduced the latter to serve as a baseline for the performance evaluation of the MPC while addressing the shortcomings of the PID control stack which is implemented in the Anafi firmware and provided by Parrot Drone.
IV-B Hardware
@@
In the real-world experiments, all ROS components were run on a laptop featuring Ubuntu 20, an Intel Core i7-6600U CPU, an Intel HD Graphics 520 GPU, 32 GB RAM and a 512 GB SSD.
IV-C Integration of the Anafi Drone into a Multi-Robot Simulation
@@
IV-D Verification of the Custom PID-based Controller Framework and of the MPC
@@
V Results
V-A Integration of the Anafi Drone into a Gazebo Simulation
@@
Since Gazebo can have difficulties loading and rendering multiple complex meshes simultaneously, we neglect the display of rotating propellers to keep the model of the Anafi drone as simple as possible. On our setup described in Section IV where both simulators run on a single computer, we achieved a real time factor of $RTF=0.6$. However, one advantage of the ROS middleware used for exchanging data between the simulators is that it can run on distributed systems, thus providing the possibility to increase the real time factor by running different components of the system illustrated by Figure 3 on different machines. As a consequence, a real time factor of $RTF=1$ was achieved when Sphinx was installed on a separate computer with near identical specifications than the computer hosting Gazebo.
V-B Evaluation of Tracking Controllers
Computational Workload
On the computational platform used for the simulation experiments, the CPU time required by our customized PID framework was approximately 2.59 times greater than that of the MPC, indicating a substantially higher computational burden. This increased cost is likely attributable to the distributed architecture of the PID framework, wherein each control cascade is deployed as an independent node. Additional nodes are responsible for tasks such as computing relative positional information between the drone and its waypoints, as well as managing the collection and redistribution of data among the various control cascades. In contrast, the MPC is formulated as a quadratic optimization problem, which can be solved with high computational efficiency.
Evaluation Scenarios
@@ V-B1 Static Waypoints
Figure 7 illustrates the same scenario but in the real world experiments. It can be seen that this time the MPC is not beneficial in comparison to the PID controller framework. However, this could be attributed to the windy conditions ($v_{wind}=4$\mathrm{m}\mathrm{/}\mathrm{s}$$ with considerably higher wind gusts possibly present when testing the MPC). Another possible explanation are the cumulated error states in combination with real-world effects not modelled in Sphinx that could lead to a slightly more pronounced overshoot of the waypoint than with the PID controller framework. However, their advantage in dynamic waypoint scenarios outweighs this aspect.
V-B2 Straight Line Trajectory
@@ V-B3 Circular Trajectory
The MPC shows better performance than the PID across all scenarios, although at higher velocities the advantage of the MPC diminishes. The reason is that in these scenarios the control commands produced by the MPC are bound by the maximum attitude angles of the drone thus limiting its manueverability. The real world experiments confirm the simulation results, although they were conducted under windy conditions subjecting the drone to heavy disturbances. However, the standard deviation of the tracking error $e_{\sigma}$ of the MPC is considerably lower than the one of the PID controller framework thus indicating that the tracking is managed in an evenly way.
V-B4 Vertical Rectilinear Periodic Trajectory
@@
The question of which tracking accuracy is required strongly depends on the mission. Whereas landing on a moving platform may only allow a tolerance within the range of less than $0.5$\mathrm{m}$$, high altitude surveillance tasks might accept a tracking error of several meters. However, the detailed experiment evaluation presented above allows to compensate the remaining tracking error by shifting the waypoint in the direction of motion based on the mean tracking error associated with the current speed of the Anafi drone.
VI Conclusion
This work integrates the Parrot Sphinx simulator with Gazebo-based multi-robot frameworks and addresses limitations of the Anafi drone’s default PID controller. We introduce two alternative controllers for tracking tasks: an improved PID controller with limited tracking accuracy, and a Model Predictive Controller (MPC) that reduces tracking error using a modified dynamics model with cumulative error states. Integration is demonstrated in our airship simulation [10], and both controllers are evaluated in simulation and real-world scenarios. We aim for this work to support the broader adoption of the Anafi drone in scientific research applications. Future work will address the current limitation of supporting only a single Anafi drone instance within a Gazebo simulation environment.